OpenStudy (anonymous):

what are the solutions for the quadratic equation x^2+6x=16

OpenStudy (swissgirl):

x^2+6x-16=0

OpenStudy (anonymous):

you can factor this quadratic.

OpenStudy (anonymous):

(x+8)(x-2)

OpenStudy (anonymous):

then set set them both equal to zero

OpenStudy (anonymous):

you get x=-8 and x=2

OpenStudy (anonymous):

so this is what i did... (x-2)(9x+8) x^2+8x-2x-16 x^2+6x-16

OpenStudy (anonymous):

what do i do after tht?

OpenStudy (anonymous):

oops! the fist step should say (x-2)(x+8)

OpenStudy (anonymous):

set them both equal to zero

OpenStudy (anonymous):

(x-2)=0 and (x+8)=0

OpenStudy (swissgirl):

did u follow?

OpenStudy (anonymous):

ohh okay yeah i got it thnks!